# Efficient Long-duration Talking Video Synthesis with Linear Diffusion Transformer under Multimodal Guidance

### v3

@@ IV-F Ablation Study
In this section, we analyze the impact of different fusion schemes as illustrated in Fig. 4 on guidance using reference portrait and audio. Specifically, for analyzing Siamese Fusion, we have to design Siamese transformers for both reference portrait and audio to produce guidance for each block of the backbone. For the reference portrait, we leverage a DiT model pre-trained on ImageNet. For the audio, we define a transformer whose blocks are equipped with two-layer MLPs. The parameters of these transformers are simultaneously updated during training.
@@ IV-G User Study
To comprehensively assess the perceptual quality of our approach, we conducted a user study with 12 participants. Each participant was shown videos generated by three different methods: Hallo, EchoMimic, and our proposed LetsTalk-L. For the evaluation, 20 audio samples were randomly selected, and for each audio, participants were asked to rank the three corresponding videos according to two criteria: realism, which measures how natural and lifelike the video appears, and synchronization, which evaluates how well the lip movements match the audio.
The results of the user study are summarized in Fig 10. The table reports the percentage of times each method was ranked first, second, or third for both realism and synchronization. As shown in the results, LetsTalk-L achieves the highest proportion of first-place rankings in both categories. These findings indicate that users consistently preferred the videos generated by our method, both in terms of visual realism and the accuracy of lip synchronization with the audio. Overall, the user study demonstrates the effectiveness of our approach in producing high-quality, realistic, and well-synchronized talking head videos.
@@ IV-H Long-Duration Generation
To validate the effectiveness of our proposed approach for long video generation, we conducted experiments to analyze the quality of 100 consecutive video clips (about 64 second) generated under three settings: (1) the baseline model (LetsTalk-L) with with the memory bank, (2) the method augmented with regularized training strategy, and (3) the method further enhanced with the patch drop augmentation. Our results in Table VII demonstrate that the memory bank significantly improves temporal consistency, while the regularized training strategy further mitigates degradation in later clips, leading to higher overall video quality. The results demonstrate that regularized training strategy primarily enhances single-frame quality, significantly reducing FID from 52.243 to 32.746, while maintaining temporal consistency (FVD 264.709). On the other hand, patch drop augmentation further improves long-term motion coherence, achieving the best FVD (251.684) with only a slight FID increase (34.821), confirming its complementary role in stabilizing long-duration generation.
@@
IV-I Visualization
We provide video visualizations showcasing long-duration results of up to 2 minutes. In these videos, each row corresponds to a different portrait driven by the same audio segment, demonstrating our model’s ability to generate consistent and diverse talking head videos across various identities. Notably, our model is trained only on HDTF [55] and CelebV-HQ [56], which feature English conversational scenarios with real portraits. To further demonstrate the generalization ability of our approach, we also present results where real portraits are driven by Chinese speech and singing audio, even though such scenarios are not included in the training data.
@@ V Conclusion
In this work, we present LetsTalk, a novel diffusion transformer framework for efficient and high-quality long-duration talking video synthesis under multimodal guidance. LetsTalk introduces a memory bank mechanism and noise-regularized training to explicitly maintain contextual continuity and mitigate error accumulation in extended video generation. Our spatiotemporal modeling, based on a deep compression autoencoder and linear attention, achieves state-of-the-art visual fidelity with substantially fewer parameters and faster inference. Through systematic analysis of multimodal fusion, we provide principled guidance for multimodal conditioning in video generation, demonstrating that deep Symbiotic Fusion for portrait features and shallow Direct Fusion for audio ensures both strong identity preservation and expressive, synchronized motion. Extensive experiments show that LetsTalk consistently outperforms existing approaches in quality, temporal coherence, and efficiency, while generalizing robustly to out-of-domain scenarios. Overall, LetsTalk advances the frontier of realistic, scalable talking video generation, paving the way for broader applications in digital humans and interactive media.

### v4

@@ IV-F Ablation Study
In this section, we analyze the impact of different fusion schemes as illustrated in Fig. 4 on guidance using reference portrait and audio.
@@ IV-G User Study
We conducted a user study with 12 participants to assess the perceptual quality of our approach. Each participant viewed videos generated by Hallo, EchoMimic, and our proposed LetsTalk-L, using 20 randomly selected audio samples. For each sample, participants ranked the three methods by realism (naturalness) and synchronization (lip-audio match). The results of the user study are summarized in Fig 8. The table reports the percentage of times each method was ranked first, second, or third for both realism and synchronization. As shown in the results, LetsTalk-L achieves the highest proportion of first-place rankings in both categories. LetsTalk-L consistently received the most first-place votes in both categories, indicating clear user preference. This demonstrates that our method produces videos with superior visual realism and lip synchronization.
@@ IV-H Long-Duration Generation
We evaluated our approach for long video generation using 100 consecutive clips ( 64 seconds) under three settings: (1) baseline LetsTalk-L with memory bank, (2) regularized training, and (3) additional patch drop augmentation. As shown in Table VII, the memory bank improves temporal consistency, regularized training boosts single-frame quality (FID drops from 52.243 to 32.746, FVD 264.709), and patch drop augmentation further enhances motion coherence, achieving the lowest FVD (251.684) with minimal FID increase (34.821). These results confirm that each technique effectively improves long-duration video quality.
@@ V Conclusion
In this work, we present LetsTalk, a novel diffusion transformer framework for efficient and high-quality long-duration talking video synthesis under multimodal guidance. LetsTalk introduces a noise-regularized memory bank to explicitly maintain contextual continuity and mitigate error accumulation in extended video generation. Our spatiotemporal modeling, based on a deep compression autoencoder and linear attention, achieves state-of-the-art visual fidelity with substantially fewer parameters and faster inference. Through systematic analysis of multimodal fusion, we provide principled guidance for multimodal conditioning in video generation, demonstrating that deep Symbiotic Fusion for portrait features and shallow Direct Fusion for audio ensures both strong identity preservation and expressive, synchronized motion. Extensive experiments show that LetsTalk consistently outperforms existing approaches in quality, temporal coherence, and efficiency, while generalizing robustly to out-of-domain scenarios. Overall, LetsTalk advances the frontier of realistic, scalable talking video generation, paving the way for broader applications in digital humans and interactive media.